Far Side Corporation is expected to pay the following dividends over the next four years: $11, $9, $5, and $2. Afterward, the company pledges to maintain a constant 8 percent growth rate in dividends forever.

  

Required:

If the required return on the stock is 16 percent, what is the current share price? (Do not round your intermediate calculations.)

Options

$33.62

$35.39

$36.45

$34.79

$43.33
Answer #1

Step 1: Computation of market price at the end of year 4 using Gordon Growth Model

P4 = D5 / (Ke – g)

Where,

P4 - Market price at the end of year 4 = ?

D5 - Expected dividend in year 5 = 2*1.08 = 2.16

Ke – Cost of equity = 16%

G – Growth rate in dividend = 8%

P4 = 2.16/(.16-.08)

= 2.16/.08

= $27.00

Step 2: Computing current share price by discounting the cashflow at required return

Year Dividend PVF@16% Present Value (Cashflow*PVF)
1 11            0.862 9.48
2 9            0.743 6.69
3 5            0.641 3.20
4 29(2+27)            0.552 16.02

current share price = Cashflow*PVF

= 9.48+6.69+3.20+16.02

= $35.39

You can use the equation 1/(1+i)^n to find PVF using calculator